OpenStudy (anonymous):

Last question hopefully. This is elimination, explain to me how I can solve. 3x +3y=27 x+3y=-11

OpenStudy (solomonzelman):

you are eliminating the variable. taking this word "eliminating" literally. \(\normalsize\color{royalblue}{ \rm 3x+3y=27 }\) \(\normalsize\color{royalblue}{ \rm x+3y=-11 }\) I would multiply the second equation times -1, and add the equations together.

OpenStudy (anonymous):

What am I multiplying by -1?

OpenStudy (solomonzelman):

every term in the second equation.

OpenStudy (solomonzelman):

what will your second equation look like?

OpenStudy (anonymous):

-1x*3y=-11 .-.

OpenStudy (solomonzelman):

\(\normalsize\color{blueviolet}{ x+3y=-11 }\) \(\normalsize\color{blueviolet}{ (-1)x+(-1)3y=(-11)(-1) }\) \(\normalsize\color{blueviolet}{ -x-3y=11 }\)

OpenStudy (solomonzelman):

you will be getting; \(\normalsize\color{teal}{ 3x +3y=27 }\) \(\normalsize\color{teal}{ -x-3y=11 \large\color{white}{|}}\) makes sense? then add the equations together

OpenStudy (anonymous):

Yes.

OpenStudy (solomonzelman):

when you add then, what do you get?

OpenStudy (anonymous):

I got 8 and 4.5

OpenStudy (solomonzelman):

for what?

OpenStudy (anonymous):

Solving both equations. I think I ended up substituting.

OpenStudy (solomonzelman):

you solved for x incorrectly whether 8 or 4.5 is for x.

OpenStudy (solomonzelman):